In a triangle,prove that sin 3A sin(B-C)+sin 3Bsin(C-A)+sin3C(A-B)=0

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Vikas TU
4 years ago
Dear student 
Dear student
Please approach using
( Sin A/ a) = ( sin B / b ) = ( sin C / c) = k
Expand sin 3A = 3ak - 4 ( ak)^3
If doubt still persists kindly revert for complete solution
Regards